use the pythagorean theorem to find the indicated quantity. Round your answer to the nearest hundredth.
10^2 = 6^2 + b^2

Answer by richwmiller(17219)


10^2 = 6^2 + b^2
100=36+b^2
64=b^2
+\-8=b